### v1

@@ 7 Barriers
Our main goal in this chapter has been to show the main principles behind the utility of the hard-core model in graph theory, especially via the local occupancy framework. Early on, in the discussion shortly following (9.2), we mentioned that the condition we use cannot be improved in general. Put in another, perhaps blatantly obvious, way, the local occupancy condition is sharp for cliques in the most general (bounded-degree) circumstances, in which neighbourhoods are permitted maximum density.
We next discuss sharpness and barriers nearer the opposite end with respect to local density – that is, triangle-free graphs. We first recap our analysis for triangle-free graphs. We established local $(\beta,\gamma)$-occupancy with specific functions $\beta$ and $\gamma$ of $\lambda$, whereby the optimization for a suitable choice of $\lambda$ tells us that $\beta+\gamma\Delta$ can be as small as $(1+o(1))\Delta/\log\Delta$ as $\Delta\to\infty$. As corollaries of Theorem 3.5, Theorem 3.6, and Theorem 5.5, we deduced, for any triangle-free $G$ of maximum degree $\Delta$, the following statements, as $\Delta\to\infty$, in increasing order of strength:
@@
The bounded-degree independence number bound, however, is more general than the Ramsey number bound, and it is matched more closely by other constructions in a regime with lower edge-density. We take $p=p(n)$ so that $d=np$ satisfies $d=\omega(\log n)$ and $d=o(n^{1/3})$, and consider the binomial random graph ${\mathbb{G}}_{n,p}$. Classic first moment estimates (see, for example, [9]) show that $\Delta({\mathbb{G}}_{n,p})\lesssim d$, that $\alpha({\mathbb{G}}_{n,p})\lesssim(2n\log d)/d$, and that ${\mathbb{G}}_{n,p}$ has $o(n)$ triangles with probability tending to $1$ as $n\to\infty$. Thus, there is some triangle-free induced subgraph $G=(V,E)$ of ${\mathbb{G}}_{n,p}$ of maximum degree $\Delta\sim d$ with $|V|\sim n$, for which $\alpha(G)\lesssim(2|V|\log\Delta)/\Delta$ with probability tending to $1$ as $n\to\infty$. This triangle-free construction shows how, over this range of densities, the triangle-free local occupancy optimization is sharp up to an asymptotic factor of at most $2$.
@@
as $\Delta\to\infty$, where $X$ is the random independent set. As the occupancy fraction is monotone increasing in terms of $\lambda$, we maintain the same occupancy fraction guarantee for $X$ being a uniformly random independent set:
@@
To illustrate this, here is an important related problem where we would be delighted to make significant improvements through such methods, but are unable to do so. Let $G$ be a bipartite graph of maximum degree $\Delta$. How large, as a function of $\Delta$, can the list chromatic number $\chi_{\ell}(G)$ of $G$ be? The rub here is that, while $G$ is necessarily triangle-free and thus has excellent local structure, the methods that we have used in this chapter do not ‘see’ (and so cannot take advantage of) the assumed global structure, the division of the vertex-set into two independent sets. Alon and Krivelevich [8] have boldly conjectured that the answer to the above question is at most $O(\log\Delta)$ – a quantity that is dramatically smaller than $\Theta(\Delta/\log\Delta)$ which we have in the triangle-free case. Unfortunately, ‘non-local’ methods have not made much headway either (see [7] and [16] for recent progress).
More subtly, let us recall two of the applications described in Section 4. Corollaries 4.7 and 4.9 both smoothly generalize the triangle-free case, and thus skirt close to the barriers, but these bounds deteriorate as they stray from the triangle-free case. More precisely, they become no better than the trivial bounds if, respectively, we forbid having $t$ edges induced in any neighbourhood for some $t=\Omega(\Delta^{2})$, or we forbid as a subgraph a cycle $C_{k}$ of length $k=\Omega(\Delta)$. Thus, the local occupancy method seems to break down with non-vanishing local edge-density. On the other hand, we still know decent bounds in this regime – that is, there is a proper colouring using at most $(1-\varepsilon)\Delta$ colours for some positive fixed $\varepsilon$, where $\varepsilon$ depends on the local edge-density bound. It might come as a surprise that a seemingly blunter method succeeds in this case: the so-called ‘naïve colouring’ method, whereby each vertex is assigned a uniformly random colour from some ground set of colours and then conflicts (that is, monochromatic edges) are resolved by a simple follow-up recolouring procedure. See [57] for a broader treatment of this method, and [43] for the most recent developments along such lines.
@@ 8 Open problems
The first problem concerns occupancy fraction. Given $\lambda>0$, let $f_{\lambda}(d)=\lambda/(1+(d+1)\lambda)$. We showed in the Introduction that, given a graph $G$ of maximum degree $\Delta$, the occupancy fraction of the hard-core model at fugacity $\lambda$ satisfies $\alpha_{G}(\lambda)\geq f_{\lambda}(\Delta)$, and that this is sharp when $G$ is a disjoint union of $K_{\Delta+1}$s. We propose the following conjecture.
@@
As $f_{\lambda}$ is convex, this summation is bounded from below by $f_{\lambda}(\overline{\deg}(G))|V|$, where $\overline{\deg}(G)$ is the average degree of $G$. Thus, Conjecture A is a stronger form of the first occupancy fraction bound we gave, while remaining sharp when $G$ is a disjoint union of complete graphs. Its proof would moreover simultaneously refine and strengthen the bound in (9.5), the Caro–Wei theorem (see [22] and [73]), and an analogous lower bound on $\log Z_{G}(\lambda)$ given by Sah, Sawhney, Stoner and Zhao [62, Theorem 1.7]. Conjecture A implies the latter result by integration (see (9.3)).
The second problem concerns the relationship between occupancy fraction and the independence number. We have already observed how, in the complete graph $K_{\Delta+1}$, the average size of an independent set is $(\Delta+1)/(\Delta+2)$, while the independence number is $1$, and so the ratio approaches $1$ as $\Delta\to\infty$. Because of the barriers discussed above, it is impossible to improve upon Shearer’s bound $R(3,k)\lesssim k^{2}/\log k$ by a naïve application of Theorem 4.2, which requires small $\lambda$ for the best quantitative bounds. But there remains a route to move further beyond local occupancy – that is, if, for all triangle-free graphs $G$, the ratio between $\alpha(G)$ and the average size of an independent set in $G$ is bounded above $1$, then one obtains an improvement to the constant factor in Shearer’s bound. The following conjecture from [27] asserts that, in triangle-free graphs, the ratio approaches $2$ as the minimum degree grows. By a simple argument that removes small degree vertices, the following conjecture if true would imply that $R(3,k)\lesssim k^{2}/(2\log k)$.
@@
Last, let $G$ be a $K_{4}$-free graph of maximum degree $\Delta$. Thus, every neighbourhood of $G$ induces a triangle-free graph. Under this constraint, there may still be non-trivial edge-density in neighbourhoods, and this falls into more difficult territory; however, a suitable local occupancy optimization (see [29]) yields in any case that there must be an independent set of size $(1+o_{\Delta}(1))(n\log\Delta)/(\Delta\log\log\Delta)$, which is currently the best bound. There is some distance to cover, as the longstanding conjecture of Ajtai, Erdős, Komlós and Szemerédi [1] posits much more.

### v2

@@ 7 Barriers
Our main goal in this chapter has been to show the main principles behind the utility of the hard-core model in graph theory, especially via the local occupancy framework. Early on, in our discussion shortly following (8.2), we mentioned that the condition we use cannot be improved in general. Put in another, perhaps blatantly obvious, way, the local occupancy condition is sharp for cliques in the most general (bounded-degree) circumstances, in which neighbourhoods are permitted maximum density.
We next discuss sharpness and barriers nearer the opposite end with respect to local density – that is, triangle-free graphs. We first recap our analysis for triangle-free graphs. We established local $(\beta,\gamma)$-occupancy with specific functions $\beta$ and $\gamma$ of $\lambda$, whereby the optimization for a suitable choice of $\lambda$ tells us that $\beta+\gamma\Delta$ can be as small as $(1+o(1))\Delta/\log\Delta$ as $\Delta\to\infty$. As corollaries of Theorem 3.3, Theorem 3.4, and Theorem 5.5, we deduced the following statements, for any triangle-free $G$ of maximum degree $\Delta$, in increasing order of strength, as $\Delta\to\infty$:
@@
The bounded-degree independence number bound, however, is more general than the Ramsey number bound, and it is matched more closely by other constructions in a regime with lower edge-density. We take $p=p(n)$ so that $d=np$ satisfies $d=\omega(\log n)$ and $d=o(n^{1/3})$, and consider the binomial random graph ${\mathbb{G}}_{n,p}$. Classic first moment estimates (see, for example, [10]) show that $\Delta({\mathbb{G}}_{n,p})\lesssim d$, that $\alpha({\mathbb{G}}_{n,p})\lesssim(2n\log d)/d$, and that ${\mathbb{G}}_{n,p}$ has $o(n)$ triangles, with probability tending to $1$ as $n\to\infty$. Thus, there is some triangle-free induced subgraph $G=(V,E)$ of ${\mathbb{G}}_{n,p}$ of maximum degree $\Delta\sim d$ with $|V|\sim n$, for which $\alpha(G)\lesssim(2|V|\log\Delta)/\Delta$ with probability tending to $1$ as $n\to\infty$. This triangle-free construction shows how, over this range of densities, the triangle-free local occupancy optimization is sharp up to an asymptotic factor of at most $2$.
@@
as $\Delta\to\infty$, where $X$ is the random independent set. As the occupancy fraction is monotone increasing in terms of $\lambda$, we maintain the same occupancy fraction guarantee when $X$ is a uniformly random independent set:
@@
To illustrate this, here is an important related problem where we would be delighted to make significant improvements through such methods, but have been unable to do so. Let $G$ be a bipartite graph of maximum degree $\Delta$. How large, as a function of $\Delta$, can the list chromatic number $\chi_{\ell}(G)$ of $G$ be? The rub here is that, while $G$ is necessarily triangle-free and thus has excellent local structure, the methods that we have used in this chapter do not ‘see’ (and so cannot take advantage of) the assumed global structure, the division of the vertex-set into two independent sets. Alon and Krivelevich [9] have boldly conjectured that the answer to the above question is at most $O(\log\Delta)$ – a quantity that is drastically smaller than $\Theta(\Delta/\log\Delta)$ which we have in the triangle-free case. Unfortunately, ‘non-local’ methods have not made much headway either (see [8] and [18] for recent progress).
More subtly, let us recall two of the applications described in Section 4. The corollaries following Theorems 4.5 and 4.6 each smoothly generalize the triangle-free case, and thus skirt close to the barriers, but these bounds deteriorate as they stray from the triangle-free case. More precisely, they become no better than the trivial bounds if, respectively, we forbid having $t$ edges induced in any neighbourhood for some $t=\Omega(\Delta^{2})$, or we forbid as a subgraph a cycle $C_{k}$ of length $k=\Omega(\Delta)$. Thus, the local occupancy method seems to break down with non-vanishing local edge-density. On the other hand, we still know decent bounds in this regime – that is, there is a proper colouring using at most $(1-\varepsilon)\Delta$ colours for some positive fixed $\varepsilon$, where $\varepsilon$ depends on the local edge-density bound. It might come as a surprise that a seemingly blunter method succeeds in this case: the so-called ‘naïve colouring’ method, whereby each vertex is assigned a uniformly random colour from some ground set of colours and then conflicts (that is, monochromatic edges) are resolved by a simple follow-up recolouring procedure. See [63] for a broader treatment of this method, and [48] for the most recent developments along such lines.
@@ 8 Open problems
The first problem concerns occupancy fraction. Given $\lambda>0$, let $f_{\lambda}(d)=\lambda/(1+(d+1)\lambda)$. We showed in Section 1 that, given a graph $G$ of maximum degree $\Delta$, the occupancy fraction of the hard-core model at fugacity $\lambda$ satisfies $\alpha_{G}(\lambda)\geq f_{\lambda}(\Delta)$, and that this is sharp when $G$ is a disjoint union of $K_{\Delta+1}$s. We propose the following conjecture.
@@
As $f_{\lambda}$ is convex, this summation is bounded from below by $f_{\lambda}(\overline{\deg}(G))|V|$, where $\overline{\deg}(G)$ is the average degree of $G$. Thus, Conjecture A is a stronger form of the first occupancy fraction bound we gave, while remaining sharp when $G$ is a disjoint union of complete graphs. Its proof would moreover simultaneously refine and strengthen the bound in (8.6), the Caro–Wei theorem (see [26] and [78]), and an analogous lower bound on $\log Z_{G}(\lambda)$ given by Sah, Sawhney, Stoner and Zhao [68, Theorem 1.7]. Conjecture A implies the latter result by integration (see (8.3)).
The second problem concerns the relationship between occupancy fraction and the independence number. We have already observed how, in the complete graph $K_{\Delta+1}$, the average size of an independent set is $(\Delta+1)/(\Delta+2)$, while the independence number is $1$, and so their ratio approaches $1$ as $\Delta\to\infty$. Because of the barriers discussed above, it is impossible to improve upon Shearer’s bound $R(3,k)\lesssim k^{2}/\log k$ by a naïve application of Theorem 4.2, which requires small $\lambda$ for the best quantitative bounds. But there remains a route to move further beyond local occupancy – that is, if, for all triangle-free graphs $G$, the ratio between $\alpha(G)$ and the average size of an independent set in $G$ is bounded above $1$, then one obtains an improvement to the constant factor in Shearer’s bound. The following conjecture from [31] asserts that, in triangle-free graphs, the ratio approaches $2$ as the minimum degree grows. By a simple argument that removes small-degree vertices, the following conjecture if true would imply that $R(3,k)\lesssim k^{2}/(2\log k)$.
@@
Last, let $G$ be a $K_{4}$-free graph of maximum degree $\Delta$, so that every neighbourhood of $G$ induces a triangle-free graph. Under this constraint, there may still be non-trivial edge-density in neighbourhoods, and this falls into more difficult territory; however, in any case a suitable local occupancy optimization (see [33]) yields an independent set of size $(1+o_{\Delta}(1))(n\log\Delta)/(\Delta\log\log\Delta)$, which is currently the best bound. There is some distance to cover, as the longstanding conjecture of Ajtai, Erdős, Komlós and Szemerédi [2] posits much more.
@@
Acknowledgements
Ewan Davies was supported in part by NSF grant CCF-2309707. Ross Kang was supported by Vidi (639.032.614) and Open Competition (OCENW.M20.009) grants of the Dutch Research Council (NWO) as well as the Gravitation programme NETWORKS (024.002.003) of the Dutch Ministry of Education, Culture and Science (OCW).
Notes added
We posted a version of this book chapter in a public preprint repository in January of 2025. Not long afterwards came a number of interesting related developments.
In March 2025 Buys, van den Heuvel and Kang [20] generalized Shearer’s proof [71] that any $n$-vertex triangle-free graph of average degree $d$ contains an independent set of size $(1-o(1))n\log(d)/d$ to the partition function of the hard-core model. This yields the lower bound that the number of independent sets (of all sizes) in such a graph must be at least
$\exp\left((1-o(1))\frac{(\log d)^{2}}{2d}n\right),$
which is best possible (keeping in mind (8.3)).
In April 2025 Klartag [59] defined a randomized construction of a large origin-symmetric ellipsoid in $\mathbb{R}^{d}$ containing no points of $\mathbb{Z}^{d}$ other than the origin. His analysis of the volume of this ellipsoid gives the existence of lattice-based sphere packings that certify the new density lower bound $\theta(d)\geq\Omega(d^{2}\cdot 2^{-d})$, improving on the result in [24].
While Shearer’s remains the state of the art for upper bounds on $R(3,k)$, in May 2025 Campos, Jenssen, Michelen and Sahasrabudhe [25] improved the lower bound on $R(3,k)$ to $(1/3-o(1))k^{2}/\log k$ as $k\to\infty$ by analyzing a variant of the triangle-free process that starts from a somewhat structured graph. It seems that starting the process carefully and steering it along a simpler trajectory is easier to analyze and yields a better bound.
Also in May 2025, Davies, Sandhu and Tan [34] showed that bounds of the form (8.5) can be used to make progress on Conjecture A. They established the conjecture in graphs of maximum degree $\Delta$ for values $\lambda\leq 3/(\Delta+1)^{2}$ (alternatively, for each $n\geq 1$ the conjecture holds in all $n$-vertex graphs for $\lambda\leq 3/n^{2}$).